OpenStudy (anonymous):

Solve the equation. 4x - 2 = -6

OpenStudy (anonymous):

-6+2=-4 4x(times) x= -4 x=-1

OpenStudy (anonymous):

\[4x-2=-6\] \[4x-2+2=-6+2\] \[4x=-4\] \[\frac{4x}{4}=\frac{-4}{4}\] \[\frac{\cancel {4}x}{\cancel{4}}=-1\] \[x=-1\]